OpenStudy (dls):

Integrate :

OpenStudy (dls):

\[\Huge \int\limits \frac{\sin x}{\sin 4x}\]

OpenStudy (zzr0ck3r):

by parts i assume

OpenStudy (dls):

didn't work

OpenStudy (dls):

sin4x=2sin2xcos2x=4sinxcosxcos2x \[\Huge \int\limits \frac{\cancel{\sin x}}{4 \cancel{\sin x} \cos x \cos 2x}= \int\limits \frac{dx}{\cos x \cos 2x}\]

OpenStudy (dls):

also 1/4

OpenStudy (dls):

\[\Huge \frac{ 1}{4} \int\limits \frac{dx}{ \cos x \cos 2x}\]

OpenStudy (zzr0ck3r):

this seems very hard by the answer they gave...

OpenStudy (zzr0ck3r):

still want it?

OpenStudy (dls):

if there's a simpler way

Miracrown (miracrown):

subsitution maybe

OpenStudy (anonymous):

\[Cos(2x) = \cos^2(x)+\sin^2(x) = 1-2\sin^2(x)=2\cos^2(x)-1\]

Miracrown (miracrown):

another way to do it would be to use a trig substitution since sin (4x) = 2 sin (2x) cos (2x)

OpenStudy (anonymous):

This one is gross

OpenStudy (anonymous):

\[\frac{ 1 }{ \sin4x } = \csc4x\] try

OpenStudy (anonymous):

sinxcsc4x = ?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sinxcsc4x \[\int\limits \frac{ 1 }{ 2\cos^3x+2cosx-6\sin^2xcosx } dx\]

OpenStudy (anonymous):

This will get you to the answer, but will take you a while...

OpenStudy (anonymous):

@Kainui need ure halp plz

OpenStudy (kainui):

Sorry I was afking

OpenStudy (kainui):

My first thought is to make the substitution 4x=u so that you're messing with the numerator instead of the denominator.

OpenStudy (kainui):

but idk lol

OpenStudy (anonymous):

Don't think that works idk

OpenStudy (kainui):

Are we allowed to use complex numbers?

OpenStudy (anonymous):

I guess so

OpenStudy (anonymous):

\[\int\limits\limits \frac{ 1 }{ 2\cos^3x+2cosx-6\sin^2xcosx } dx \times \frac{ cosx }{ cosx }\]

OpenStudy (anonymous):

U sub then partial fractions...omg im out

OpenStudy (kainui):

@DLS are there bounds on this integral? It might be easy to solve by making symmetry arguments.

OpenStudy (dls):

nope no bounds , the next question however has bounds from 0 to 1

OpenStudy (anonymous):

Let \[I = \int \frac{sin _\ x}{sin _\ 4x} dx \] So \[ I = \int \frac{cosx _\ dx}{4cos^2 x _\ cos_\ 2x}\] \[I = \int \frac{cos_\ x _\ dx}{4(1-sin^2 x)(1-2sin^2 x)}\] Now substitute sin x = u so that \[I = \int \frac{du}{4(1-u^2 )(1-2u^2)}\] Now use partial fractions to integrate with respect to u and then inverse substitution, u = sin x. \[I = \frac{1}{4}\int \left( \frac{2}{(1-2u^2)} - \frac{1}{(1-u^2)} \right) du \] \[I = \frac{1}{8}\left( 2\ln \frac{\sqrt {2} u + 1}{(\sqrt{2}u - 1 )} -\ln \frac{1+u}{(1-u)} \right) + C\] \[I = \frac{1}{8}\left( 2\ln \frac{\sqrt {2} sin_\ x+ 1}{\sqrt{2}sin_\ x- 1 } -\ln \frac{1+sin_\ x}{1-sin_\ x} \right) + C\] Hope this helps. Alternatively, You can use complex numbers and De Moivre's theorem. You can use \[ sin _\ x = \frac{\exp (i x) - \exp(-i x)}{2i} \] and \[ sin _\ 4x = \frac{\exp (4i x) - \exp(-4i x)}{2i} \]